If Y=aX+b, show that Y has the same coefficient of skewness?


1
Expert's answer
2021-04-14T01:48:52-0400

The skewness of a random variable X is the third standardized moment, defined as "\tilde{\mu}^3(X)=\sigma(X)^{-3}E[(X-\mu(X))^3]" , where "\mu(X),\ \sigma(X)" are the mean and the standard deviation of X. If "Y=aX+b", then

"\mu(Y)=E(aX+b)=a\mu(X)+b"

"Var(Y)=E((Y-\mu(Y))^2)=E((aX+b-(a\mu(X)+b))^2)=a^2E((X-\mu(X))^2)=a^2Var(X)"

"\sigma(Y)=\sqrt{Var(Y)}=\sqrt{a^2Var(X)}=|a|\sigma(X)"

"\tilde{\mu}^3(Y)=\sigma(Y)^{-3}E[(Y-\mu(Y))^3]=|a|^{-3}\sigma(X)^{-3}E[(aX+b-(a\mu(X)+b))^3]=\frac{a^3}{|a|^3}E((X-\mu(X))^3)=\mathrm{sign}(a) \tilde{\mu}^3(X)"

Therefore, if a>0, then Y has the same coefficient of skewness as X. If a<0 then X, Y have the same absolute values of the coefficient of skewness, but their signs are different.

If a=0, then the coefficient of skewness of Y is undefined.